LASER INDUCED FLUORESCENCE STUDY OF B̃ - Ã TRANSITION OF ISOPROPOXY RADICAL

\underline{RABI CHHANTYAL-PUN}, JINJUN LIU^{a} and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus OH 43210.

Alkoxy radicals are important intermediates in combustion and atmospheric chemistry. Alkoxy radicals are also of significant spectroscopic interest for the study of Jahn-Teller and pseudo-Jahn-Teller effects, involving the X̃ and Ã states. Isopropoxy (CH_{3}CH(O)CH_{3}) radical is the simplest secondary alkoxy radical. High resolution laser induced fluorescence (LIF) studies of the B̃ - X̃ transition have been performed previously.^{b} In this talk we will present results obtained from a moderate resolution LIF study of the B̃ - Ã transition whose analysis should complement the B̃ - X̃ analysis. The separation between the Ã and X̃ band origin was found to be 58(±3) cm^{−1} which is more precise but consistent with the previous dispersed fluorescence experiment.^{c} We are also able to observe transitions corresponding to CO stretch progressions and a low frequency fundamental vibrational mode. The origin band and CO stretch progressions were found to have similar rotational contours which differ from the B̃ - X̃ origin band contour. Analysis of the different band contours of B̃ - Ã transitions and their implications will be presented.

aPresent address: Department of Chemistry, University of Louisville, 2320 South Brook Street, Louisville, Kentucky 40292

bJ. Liu and T. A. Miller, 65^{th} OSU International Symposium on Molecular Spectroscopy, The Ohio State University, Columbus, Ohio, 2010, TG-14

cJ. Jin, I. Sioutis, G. Tarczay, S. Gopalakrishnan, A. Bezant, and T. A. Miller, J. Chem. Phys. 121, 11780 (2004)